# Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain

### v1

@@ 4.2 Communicating Structures
If the adversary is communicating, $\mathcal{P}_{s},s\in S$ are convex and compact, and $\mathcal{Q}$ is convex, then the constant-gain average-reward robust Bellman equation (3.1) has a solution.
5 Bellman Optimality for the HD-S Case
In this section, we show a surprising result that, under a weak communication assumption, the average reward for a history-dependent controller against a stationary adversary corresponds to the solution of the Bellman equation with the inf-sup ordering, rather than its original form.
Definition 4.
A stationary adversary policy $p\in\mathcal{P}$ is said to be weakly communicating if there is a communicating class $C$ s.t. for any $s,s^{\prime}\in C$, there exists $\Delta:S\rightarrow\mathcal{Q}$ and $N\geq 1$ s.t. $p_{\Delta}^{N}(s^{\prime}|s)>0$. Moreover, for all $\in C^{c}, ss$ is transient under any stationary policy of the controller. The adversary is weakly communicating if every stationary policy $p\in\mathcal{P}$ is weakly communicating. Evidently, if the adversary is communicating, then it is also weakly communicating.
Proposition 5.1.
If $\left\{{\delta_{a}:a\in A}\right\}\subset\mathcal{Q}$ and the adversary is weakly communicating, then for each $\epsilon>0$, there exists a history-dependent reinforcement learning policy $\pi_{\mathrm{RL}}\in\Pi_{\mathrm{H}}$ s.t.
$0\leq\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)\leq\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\sup_{\pi\in\Pi_{\mathrm{S}}}\underline{\alpha}(\mu,\pi,\kappa)-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)\leq\epsilon$
Theorem 6.
If $\left\{{\delta_{a}:a\in A}\right\}\subset\mathcal{Q}$, then so long as the adversary is weakly communicating,
@@ 5 Bellman Optimality for the HD-S Case
The same result holds true if $\underline{\alpha}$ is replaced with $\overline{\alpha}$. Moreover, if (3.2) admits a solution pair $(u^{\prime},\alpha^{\prime})$, then $\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})\;=\;\overline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})\;=\;\alpha^{\prime}.$ A sufficient condition for (3.2) to have a solution is that the adversary is communicating and $\mathcal{P}_{s}:s\in S$ are compact.
@@
Assume that both the controller and the adversary are communicating, $\left\{{\delta_{a}:a\in A}\right\}\subset\mathcal{Q}$, and that $\mathcal{Q}$ and $\mathcal{P}_{s},\;s\in S$ are compact. If the adversary’s policy class is stationary, i.e., $\mathrm{K}=\mathrm{K}_{\mathrm{S}}$, then stationary policies are optimal for a history-dependent controller $\Pi=\Pi_{\mathrm{H}}$ if and only if $\alpha^{\prime}=\alpha^{*}$.
@@ A.3 Proof of Theorem 2
Appendix B Proofs for Section 4
B.1 Proof of Theorem 3
@@
Since $\overline{\mathcal{P}}_{s}$ is compact and does not depend on $\gamma,\phi,u,\alpha$, by Berge’s maximum theorem [2, VI.3, Theorem 1 & 2], the mapping
@@
B.2 Proof of Theorem 4
@@
Step 1: We will construct a finite subset of controller’s policy $C$ that ensures a uniform communication probability lower bound. We use this property to show an enhanced version of the communicating controller under compact $\mathcal{Q}$ in Lemma 1.
We consider an arbitrary fixed pair of states $x,y\in S$. Since the controller is assumed to be communicating, for any stationary policy $\Delta:S\rightarrow\mathcal{Q}$, there exists $p=p^{x,y,\Delta}\in\mathcal{P}$ and $N=N^{x,y,\Delta}\geq 1$ s.t. $p_{\Delta}^{N}(y|x)>0$. Note that both $p$ and $N$ are dependent on $\Delta,x,y$. We will suppress the dependence for notation simplicity.
Note that if we fix $p=p^{\Delta}$ and $N=N^{\Delta}$, then the mapping $\eta\rightarrow p^{N}_{\eta}$ for $\eta:S\rightarrow\mathcal{P}(\mathcal{A})$ is continuous in $\eta$. This is because both $\eta\rightarrow p_{\eta}$ and $p_{\eta}\rightarrow p_{\eta}^{N}$ are continuous. So, there must exists an open neighborhood $G_{\Delta}$ of $\Delta$ s.t. $p_{\eta}^{N}(y|x)\geq p_{\Delta}^{N}(y|x)/2$ for all $\eta\in G_{\Delta}.$
Since $\mathcal{Q}$ is compact, $\left\{{S\rightarrow\mathcal{Q}}\right\}$, seen as stochastic matrices in $\mathbb{R}^{|S|\times|A|}$, is also compact. Note that clearly, $\left\{{G_{\Delta}:\Delta\in\left\{{S\rightarrow\mathcal{Q}}\right\}}\right\}$ is an open cover of $\left\{{S\rightarrow\mathcal{Q}}\right\}$. Hence, by the compactness of $\left\{{S\rightarrow\mathcal{Q}}\right\}$, there exists a finite sub-cover $\left\{{G_{\Delta}:\Delta\in C}\right\}$ where $C:=\left\{{\Delta_{1},\dots,\Delta_{|C|}}\right\}\subset\left\{{S\rightarrow\mathcal{Q}}\right\}$ is a finite subset.
Since $C$ is finite, by the construction of $G_{\Delta}$, we have that
$\min_{\Delta\in C}(p_{\Delta}^{\Delta})^{N^{\Delta}}(y|x)=:\delta(y|x)>0.$
Note that $\delta(y|x)$ is independent of $\Delta$, and here the dependence of $N$ and $p$ on $x,y$ are suppressed. On the other hand, for any stationary policy $\eta$, $\eta\in G_{\Delta_{k}}$ for some $k\in\left\{{1,\dots,|C|}\right\}$. So,
$(p_{\eta}^{\Delta_{k}})^{N^{\Delta_{k}}}(y|x)\geq\frac{1}{2}(p_{\Delta_{k}}^{\Delta_{k}})^{N^{\Delta_{k}}}(y|x)\geq\delta(y|x)>0.$
Since the state space is finite, we can define
$\delta:=\min_{x,y\in S}\delta(y|x),\qquad M:=\max_{k=1,\dots,|C|;\;x,y\in S}N^{x,y,\Delta_{k}}.$
This implies the following Lemma 1.
Lemma 1.
Under the assumptions of Theorem 4, there exists $\delta>0$ and positive integer $M$ s.t. for any stationary controller policy $\Delta:S\rightarrow\mathcal{Q}$ and any pair of states $x,y\in S$, there exists $p\in\mathcal{P}$ and $N\leq M$ s.t. $p^{N}_{\Delta}(y|x)\geq\delta$.
Step 2: We show that under Lemma 1, for fixed $y\in S$ and policy $\Delta:S\rightarrow\mathcal{Q}$, there exists a $p\in\mathcal{P}$ independent of the initial state $x\in S$ so that the probability of the first hitting time $\tau_{y}:=\inf\left\{{k\geq 0:X_{k}=y}\right\}$ being less than or equal to $M$ has a uniform lower bound. This is summarized in Lemma 2.
Lemma 2.
Under the assumptions of Theorem 4, there exists $\delta^{\prime}>0$ s.t. for any stationary controller policy $\Delta:S\rightarrow\mathcal{Q}$ and states $y\in S$, there exists $q\in\mathcal{P}$ s.t.
$\min_{x\in S}P_{x}^{\Delta,q}(\tau_{y}\leq|S|)\geq\delta^{\prime}.$
We fix $y\in S$ and $\Delta:S\rightarrow\mathcal{Q}$. To verify Lemma 2, note that in Lemma 1 we can take $M\leq|S|$, since the longest loop-free path between two states in $S$ has at most $|S|$ steps.
For each $x\in S$, let $p=p^{x,y,\Delta}$ the adversarial stationary policy in Lemma 1. Then, we have that
$P^{\Delta,p}_{x}(\tau_{y}\leq M)\geq p_{\Delta}^{N}(y|x)\geq\delta.$
So, to prove Lemma 2, we can show that the choice $p^{x,y,\Delta}$ can be made independent of $x$.
@@
Let us define $q=q^{y,\Delta}\in\mathcal{P}$ algorithmically as follows. We will iteratively assign $q(\cdot|s,\cdot)\in\mathcal{P}_{s}$ until all $\left\{{q(\cdot|s,\cdot):s\in S}\right\}$ has been assigned. Denote $q_{\Delta}(\cdot|s)=\sum_{a\in A}\Delta(a|s_{i})q(\cdot|s,a)$.
@@
We initialize the algorithm by assigning $q(\cdot|y,\cdot)=p_{y}$ for an arbitrary $p_{y}\in\mathcal{P}_{y}$. Then, Let $V=\left\{{y}\right\}$ be the assigned states, and $V^{c}$ the compliment in $S$ are the unassigned states.
@@
    $p_{\Delta}(s_{i+1}|s_{i})\geq\prod_{i=0}^{k-1}p_{\Delta}(s_{i+1}|s_{i})\geq\delta|S|^{-N}$
    where $\delta|S|^{-N}$ is independent of $x,y,\Delta$.
@@
    Let $j=\min\left\{{i\geq 1:s_{i}\in V}\right\}$ be the first index so that $s_{j}$ is assigned. So, $s_{i}\in V^{c}$ for all $i\leq j-1$. We assign $q(\cdot|s_{i},\cdot):=p^{s_{0},y,\Delta}(\cdot|s_{i},\cdot)\in\mathcal{P}_{s_{i}}$, which implies that $q_{\Delta}(s_{i+1}|s_{i})\geq\delta|S|^{-N}$ for all $i\leq j-1$.
    Note that, at the current iteration, $q_{\Delta}(\cdot|s)$ is well-defined for all $s\in V$. Since $s_{j}\in V$, there is a non-repeating path $\left\{{s_{j}=s_{j}^{\prime},s^{\prime}_{j+1},\dots,s^{\prime}_{k}=y}\right\}\subset V$ s.t. $q_{\Delta}(s^{\prime}_{i+1}|s^{\prime}_{i})\geq\delta|S|^{-N}$.
@@
Note that the algorithm terminates in at most $|S|$ iterations, producing $q\in\mathcal{P}$. Moreover, it produces a directed graph whose edges correspond to a positive transition probabilities at least $\delta|S|^{-N}$ under $q_{\Delta}(\cdot|\cdot)$, ensuring that every state can reach $y$ in at most $|S|$ steps.
Therefore, we conclude that for any $y\in S$ and $\Delta:S\rightarrow\mathcal{Q}$, let $q=q^{y,\Delta}\in\mathcal{P}$ be constructed by the above algorithm, then
$\min_{x\in S}P_{x}^{\Delta,q}(\tau_{y}\leq|S|)\geq\left(\delta|S|^{-N}\right)^{|S|}=:\delta^{\prime}>0.$
Note that $\delta^{\prime}$ is independent of $x,y,\Delta$. This shows Lemma 2.
Step 3: We turn the probability bound in Lemma 2 to a bound on the expected hitting time bound.
Lemma 3.
Under the assumptions of Theorem 4, for any $\Delta:S\rightarrow\mathcal{Q}$ and $y\in S$, the $q\in\mathcal{P}$ in Lemma 2 satisfies
$\max_{x\in S}E^{\Delta,q}_{x}\tau_{y}\leq\frac{|S|}{\delta^{\prime}}.$
To show Lemma 3, we first show that
$\max_{x\in S}P_{x}^{\Delta,q}(\tau_{y}>m|S|)\leq(1-\delta^{\prime})^{m}.$ (B.3)
We prove by an induction on $m$. The base case $m=1$ follows directly from Lemma 2 that
$\max_{x\in S}P_{x}^{\Delta,q}(\tau_{y}>|S|)=1-\min_{x\in S}P_{x}^{\Delta,q}(\tau_{y}\leq|S|)\leq 1-\delta^{\prime}.$ (B.4)
@@
$P_{x}^{\Delta,q}(\tau_{y}>(k+1)|S|)$ $=E_{x}^{\Delta,q}\mathds{1}\left\{{\tau_{y}>(k+1)|S|}\right\}$
$=E_{x}^{\Delta,q}E_{x}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>(k+1)|S|}\right\}\middle|\mathcal{H}_{k|S|}\right]$
$\stackrel{{\scriptstyle(i)}}{{=}}E_{x}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}E_{x}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>(k+1)|S|}\right\}\middle|\mathcal{H}_{k|S|}\right]\right]$
$\stackrel{{\scriptstyle(ii)}}{{=}}E_{x}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}E_{X_{k|S|}}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>|S|}\right\}\right]\right]$
$=E_{x}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}P_{X_{k|S|}}^{\Delta,q}(\tau_{y}>|S|)\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}E_{x}^{\Delta,q}[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}](1-\delta^{\prime})$
@@
We then prove Lemma 3 using (B.3). Note that since $\tau_{y}$ is non-negative, for $x\in S$, $s\neq y$,
$E_{x}^{\Delta,q}[\tau_{y}]$ $=\sum_{k\geq 0}P_{x}^{\Delta,q}(\tau_{y}\geq k)$
$=\sum_{k\geq 0}P_{x}^{\Delta,q}(\tau_{y}>k)$
$\leq|S|+\sum_{k\geq 1}|S|P_{x}^{\Delta,q}(\tau_{y}\geq k|S|)$

### v2

@@ 6 Bellman Optimality for the HD-S Case
If $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, then so long as the adversary is weakly communicating,
@@
The same result holds true if $\underline{\alpha}$ is replaced with $\overline{\alpha}$. Moreover, if (3.2) admits a solution pair $(u^{\prime},\alpha^{\prime})$, then
$\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})\;=\;\overline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})\;=\;\alpha^{\prime}.$
@@
Assume that both the controller and the adversary are weakly communicating, $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, and that $\mathcal{Q}$ and $\mathcal{P}_{s},\;s\in S$ are compact. If the adversary’s policy class is stationary, i.e., $\mathrm{K}=\mathrm{K}_{\mathrm{S}}$, then stationary policies are optimal for a history-dependent controller $\Pi=\Pi_{\mathrm{H}}$ if and only if $\alpha^{\prime}=\alpha^{*}$.
@@ A.2 Proof of Theorem 1
$\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}}\liminf_{n\rightarrow\infty}E_{\mu}^{\Delta,\kappa}\frac{1}{n}\sum_{k=0}^{n-1}r(X_{k},A_{k})\geq\alpha^{*}-\epsilon.$ (A.4)
@@ A.3 Proof of Theorem 2
A.4 Proof of Theorem 3
Proof.
We observe that Step 1 in the proof of Theorem 1 shows that the policy $\Delta$ assumed in the statement of Theorem 3 satisfies (A.4); i.e.
By Theorem 1 and set inclusion arguments, we derive that
$\alpha^{*}=\overline{\alpha}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}})=\overline{\alpha}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}})=\underline{\alpha}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}})\geq\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\Delta,\kappa)\geq\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}}\underline{\alpha}(\mu,\Delta,\kappa)\geq\alpha^{*}-\epsilon.$
Rearranging these inequalities gives Theorem 3.
Appendix B Proof of Theorem 4
@@
Since $\overline{\mathcal{P}}_{s}$ is compact and does not depend on $\gamma,\phi,u,\alpha$, by Berge’s maximum theorem [3, VI.3, Theorem 1 & 2], the mapping
@@
Appendix C Proof of Auxiliary Lemmas in Section 5
C.1 Proof of Lemma 1
@@
Fix $\Delta$ and its communicating set $C_{\Delta}\subseteq S$; write $C_{\Delta}^{c}:=S\setminus C_{\Delta}$. We separately consider $w\in C_{\Delta}$ and $w\in C_{\Delta}^{c}$.
Case 1: $w\in C_{\Delta}$. Since $w\in C_{\Delta}$, by weak communication, there exist $p=p^{w,y,\Delta}\in\mathcal{P}$ and $N_{1}\leq|C_{\Delta}|$ such that $p_{\Delta}^{N_{1}}(y|w)>0$. This proves the claim.
Case 2: $w\in C_{\Delta}^{c}$. We choose an arbitrary $p\in\mathcal{P}$. By weak communication, there exists a non-repeating path $w=s_{0}\rightarrow s_{1}\rightarrow\cdots\rightarrow s_{k}=x\in C_{\Delta}$ with $k\leq|C_{\Delta}^{c}|$ and $\prod_{j=0}^{k-1}p_{\Delta}(s_{j+1}|s_{j})>0$. Define
$q(\cdot|s,\cdot):=p(\cdot|s,\cdot)\in\mathcal{P}_{s},\quad\forall s\in C_{\Delta}^{c}.$
Inside $C_{\Delta}$, we apply the $p^{x,y,\Delta}\in\mathcal{P}$ and $N_{1}\leq|C_{\Delta}|$ in Case 1, with $p_{\Delta}^{N_{1}}(y|x)>0$, and set
$q(\cdot|s,\cdot):=p^{x,y,\Delta}(\cdot|s,\cdot)\in\mathcal{P}_{s},\quad\forall s\in C_{\Delta}.$
By construction, under $q_{\Delta}$, there is a non-repeating path $w=s_{0}\rightarrow\cdots\rightarrow s_{k}=x=c_{0}\rightarrow\dots\rightarrow c_{N_{1}}=y$ such that $\prod_{j=0}^{k-1}q_{\Delta}(s_{j+1}|s_{j})>0$ and $k\leq|C_{\Delta}^{c}|$. Hence, $q_{\Delta}^{N}(y|w)>0$ with $N:=k+N_{1}\leq|C_{\Delta}^{c}|+|C_{\Delta}|=|S|$. Moreover, By S-rectangularity, $q\in\mathcal{P}$. This proves the claim with $p=q$. ∎
C.2 Proof of Lemma 2
Proof.
By weak communication, for each $p\in\mathcal{P}$, every state in $C_{\Delta}^{c}$ is transient under controller $\Delta$. Thus $M_{\Delta,p}^{\Delta}$ is the transient block of $p_{\Delta}$ in the canonical classification. So, letting $T_{C_{\Delta}}:=\inf\left\{{t\geq 0:X_{t}\in C_{\Delta}}\right\}$, a first transition analysis argument suggests that $I-M_{\Delta,p}^{\Delta}$ is invertible and for any $z\in C_{\Delta}^{c}$,
$(I-M_{\Delta,p}^{\Delta})^{-1}e](z)=E^{\Delta,p}_{z}T_{C_{\Delta}}<\infty.$
Consider the mapping
$g:(\eta,p)\rightarrow\det\left(I-M_{\eta,p}^{\Delta}\right).$
The mapping $(\eta,p)\rightarrow M_{\eta,p}^{\Delta}$ is continuous (entrywise), and the determinant is a polynomial in entries, hence $g$ is continuous on $\{S\to\mathcal{Q}\}\times\mathcal{P}$.
Since $\mathcal{P}$ is compact and does not depend on $\eta$, by Berge’s maximum theorem [3, VI.3, Theorem 1 & 2]
$h:\eta\rightarrow\min_{p\in\mathcal{P}}|g(\eta,p)|$
is also continuous in $\eta$.
On the other hand, at $\eta=\Delta$, one has that for every $p$, $g(\Delta,p)\neq 0$, by the invertibility of $I-M_{\Delta,p}^{\Delta}$. So, for some $p^{\prime}\in\mathcal{P}$,
$h(\Delta)=\min_{p\in\mathcal{P}}|g(\eta,p)|=\left|\det\left(I-M_{\Delta,p^{\prime}}^{\Delta}\right)\right|>0.$
Therefore, by the continuity of $h$ there exists an open neighborhood $K^{\prime}_{\Delta}$ s.t. for any $\eta\in K^{\prime}_{\Delta}$, $M_{\eta,p}^{\Delta}$ is not all 0 for any $p$, and
$h(\eta)=\min_{p\in\mathcal{P}}\left|\det\left(I-M_{\eta,p}^{\Delta}\right)\right|>0.$
Therefore, we conclude that $\forall(\eta,p)\in K^{\prime}_{\Delta}\times\mathcal{P}$, $I-M_{\eta,p}^{\Delta}$ is invertible.
With this, we define
$\phi(\eta,p):=e^{\top}(I-M_{\eta,p}^{\Delta})^{-1}e$
on $K^{\prime}_{\Delta}\times\mathcal{P}$, which is also continuous by the continuity of matrix inversion. So, applying Berge’s maximum theorem again, we conclude that by the compactness of $\mathcal{P}$,
$\eta\rightarrow\max_{p\in\mathcal{P}}\phi(\eta,p),\quad\eta\in K^{\prime}_{\Delta}.$
is continuous. Moreover, note that by the Neumann series representation, for all $(\eta,p)\in K^{\prime}_{\Delta}\times\mathcal{P}$
$e^{\top}(I-M_{\eta,p}^{\Delta})^{-1}e=\sum_{n=0}^{\infty}e^{\top}(M_{\eta,p}^{\Delta})^{n}e\geq 0$
Finally, to conclude the lemma, we note that
$0\leq\max_{p\in\mathcal{P}}\phi(\Delta,p)=\max_{p\in\mathcal{P}}e^{\top}(I-M_{\Delta,p}^{\Delta})^{-1}e<\infty,$
and by continuity, there exists an open neighborhood $K_{\Delta}\subseteq K^{\prime}_{\Delta}$ such that for all $\eta\in\overline{K_{\Delta}}$,
$0\leq\max_{p\in\mathcal{P}}e^{\top}(I-M_{\eta,p}^{\Delta})^{-1}e=\max_{p\in\mathcal{P}}\phi(\eta,p)\leq\max_{p\in\mathcal{P}}\phi(\Delta,p)+1=\max_{p\in\mathcal{P}}e^{\top}(I-M_{\Delta,p}^{\Delta})^{-1}e+1<\infty.$
This concludes the proof of Lemma 2. ∎
C.3 Proof of Lemma 4
Proof.
Consider any $w\in S$. By Lemma 3, for any $\Delta:S\rightarrow\mathcal{Q}$ with $\Delta\in G_{\Delta_{k}}$ and $y\in C_{\Delta_{k}}$, there is $p\in\mathcal{P}$ and $N\leq|S|$ (both can be dependent on $w,y,\Delta$) such that
$P^{\Delta,p}_{w}(\tau_{y}\leq|S|)\geq p_{\Delta}^{N}(y|w)\geq\delta.$
We will first show that for some $\delta^{\prime}>0$ independent of $\Delta$ and $y$, there is $q\in\mathcal{P}$ s.t.
$\min_{w\in S}P_{w}^{\Delta,q}(\tau_{y}\leq|S|)\geq\delta^{\prime};$ (C.1)
i.e. choice $p\in\mathcal{P}$ in Lemma 3 can be made independent of $w$.
@@ C.3 Proof of Lemma 4
To this end, let us define $q=q^{y,\Delta}\in\mathcal{P}$ algorithmically as follows. We will iteratively assign $q(\cdot|s,\cdot)\in\mathcal{P}_{s}$ until all $\left\{{q(\cdot|s,\cdot):s\in S}\right\}$ has been assigned.
@@
We initialize the algorithm by assigning $q(\cdot|y,\cdot)=p_{y}$ for an arbitrary $p_{y}\in\mathcal{P}_{y}$. Then, let $V=\left\{{y}\right\}$ be the assigned states, and $V^{c}=S\setminus V$ the complement in $S$ be the unassigned states.
@@
    $p_{\Delta}(s_{i+1}|s_{i})\geq\prod_{i=0}^{k-1}p_{\Delta}(s_{i+1}|s_{i})\geq\delta|S|^{-N}.$
@@
    Let $j=\min\left\{{i\geq 1:s_{i}\in V}\right\}$ be the first index so that $s_{j}$ is assigned. So, $s_{i}\in V^{c}$ for all $i\leq j-1$. We assign $q(\cdot|s_{i},\cdot):=p^{s_{0},y,\Delta}(\cdot|s_{i},\cdot)\in\mathcal{P}_{s_{i}}$. Note that the constuction of this path implies that $q_{\Delta}(s_{i+1}|s_{i})\geq\delta|S|^{-N}$ for all $i\leq j-1$.
    Moreover, at the current iteration, $q_{\Delta}(\cdot|s)$ is well-defined for all $s\in V$. Since $s_{j}\in V$, there is a non-repeating path $\left\{{s_{j}=s_{j}^{\prime},s^{\prime}_{j+1},\dots,s^{\prime}_{k}=y}\right\}\subseteq V$ s.t. $q_{\Delta}(s^{\prime}_{i+1}|s^{\prime}_{i})\geq\delta|S|^{-N}$.
@@
Note that the algorithm terminates in at most $|S|$ iterations, producing $q\in\mathcal{P}$, as we always assign $q(\cdot|s,\cdot)\in\mathcal{P}_{s}$. Moreover, it produces a directed graph whose edges correspond to a positive transition probability at least $\delta|S|^{-N}$ under $q_{\Delta}(\cdot|\cdot)$, ensuring that every state can reach $y$ in at most $|S|$ steps.
Therefore, we conclude that with $q=q^{y,\Delta}\in\mathcal{P}$ constructed by the above algorithm,
$\min_{w\in S}P_{w}^{\Delta,q}(\tau_{y}\leq|S|)\geq\left(\delta|S|^{-N}\right)^{|S|}=:\delta^{\prime}>0.$
Note that $\delta^{\prime}$ is independent of $\Delta$ and $y$. This shows (C.1).
Under a standard renewal-type argument, we turn the probability bound in (C.1) into the expected hitting time bound in Lemma 4.
First, we show that
$\max_{w\in S}P_{w}^{\Delta,q}(\tau_{y}>m|S|)\leq(1-\delta^{\prime})^{m}.$ (C.2)
We prove this by induction on $m$. The base case $m=1$ follows directly from (C.1) that
$\max_{w\in S}P_{w}^{\Delta,q}(\tau_{y}>|S|)=1-\min_{w\in S}P_{w}^{\Delta,q}(\tau_{y}\leq|S|)\leq 1-\delta^{\prime}.$ (C.3)
@@
$P_{w}^{\Delta,q}(\tau_{y}>(k+1)|S|)$ $=E_{w}^{\Delta,q}\mathds{1}\left\{{\tau_{y}>(k+1)|S|}\right\}$
$=E_{w}^{\Delta,q}E_{w}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>(k+1)|S|}\right\}\middle|\mathcal{H}_{k|S|}\right]$
$\stackrel{{\scriptstyle(i)}}{{=}}E_{w}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}E_{w}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>(k+1)|S|}\right\}\middle|\mathcal{H}_{k|S|}\right]\right]$
$\stackrel{{\scriptstyle(ii)}}{{=}}E_{w}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}E_{X_{k|S|}}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>|S|}\right\}\right]\right]$
$=E_{w}^{\Delta,q}\left[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}P_{X_{k|S|}}^{\Delta,q}(\tau_{y}>|S|)\right]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}E_{w}^{\Delta,q}[\mathds{1}\left\{{\tau_{y}>k|S|}\right\}](1-\delta^{\prime})$
@@
We then prove Lemma 4 using (C.2). Note that since $\tau_{y}$ is non-negative, for $w\in S$, $w\neq y$,
$E_{w}^{\Delta,q}[\tau_{y}]$ $=\sum_{k\geq 0}P_{w}^{\Delta,q}(\tau_{y}\geq k)$
$=\sum_{k\geq 0}P_{w}^{\Delta,q}(\tau_{y}>k)$
$\leq|S|+\sum_{k\geq 1}|S|P_{w}^{\Delta,q}(\tau_{y}\geq k|S|)$